You have found the following stock quote for RJW Enterprises, Inc., in the financial pages of todays newspaper. NET 52-WEEKIf any work can be done on a BA 2 Plus calculator please include that.
Answer #1
a. Price = Dividend/yield
=2.2/0.025=88
b. EPS = Price/(P/E) =88/15=5.86666667
Net income = EPS*shares outstanding =5.86666667*25000000=146666666.75